Cristina puts $500 in a bank account. Each year the account earns 6% simple interest. How much interest will be earned in three years?
Answer by mananth(16946)

Cristina puts $500 in a bank account. Each year the account earns 6% simple interest. How much interest will be earned in three years?
..
i= p*r*t
500
r=6%
t=3
i = 500*0.06*3
i=90
..
A= p+i
A= 500+90 = 590
she will have $590 at the end of three years
...
m.ananth@hotmail.ca